OpenStudy (anonymous):

Which is a factor of 12x2– 46x – 8 ? Answer 2 3x – 4 2x + 1 3x – 2

myininaya (myininaya):

12x^2-46x-8=2(6x^2-23x-4)=2(6x^2-24x+x-4)=2(6x(x-4)+(x-4)) =2(6x+1)(x-4)

myininaya (myininaya):

2 is a factor 6x+1 is a factor x-4 is a factor

myininaya (myininaya):

the only one listed above is 2

OpenStudy (anonymous):

thxs

OpenStudy (anonymous):

12x2– 46x – 8 = 12x2– 48x + 2x– 8 = [12x2– 48x] + [2x– 8 ] = 12x[1x– 4] + 2[2– 4 ] factors are [12x - 2] and [x + 4]

myininaya (myininaya):

you can factor one of you factors gian

OpenStudy (anonymous):

Myin you are right!

myininaya (myininaya):

did you switch your signs? 12x+2 and x-4

OpenStudy (anonymous):

Let me see..

OpenStudy (anonymous):

12x2– 46x – 8 = 12x2– 48x + 2x– 8 = [12x2– 48x] + [2x– 8 ] = 12x[x– 4] + 2[x– 4 ] = [12x+2][x-4] = 2[6x+1][x-4] so yes 2 is a factor!

myininaya (myininaya):

gj i will give u a medal gian

OpenStudy (anonymous):

You are very kind!